### v1

# Enhancing Socially-Aware Robot Navigation through Bidirectional Natural Language Conversation

Abstract
Robot navigation is an important research field with applications in various domains. However, traditional approaches often prioritize efficiency and obstacle avoidance, neglecting a nuanced understanding of human behavior or intent in shared spaces. With the rise of service robots, there’s an increasing emphasis on endowing robots with the capability to navigate and interact in complex real-world environments. Socially aware navigation has recently become a key research area. However, existing work either predicts pedestrian movements or simply emits alert signals to pedestrians, falling short of facilitating genuine interactions between humans and robots. In this paper, we introduce the Hybrid Soft Actor-Critic with Large Language Model (HSAC-LLM), an innovative model designed for socially-aware navigation in robots. This model seamlessly integrates deep reinforcement learning with large language models, enabling it to predict both continuous and discrete actions for navigation. Notably, HSAC-LLM facilitates bidirectional interaction based on natural language with pedestrian models. When a potential collision with pedestrians is detected, the robot can initiate or respond to communications with pedestrians, obtaining and executing subsequent avoidance strategies. Experimental results in 2D simulation, the Gazebo environment, and the real-world environment demonstrate that HSAC-LLM not only efficiently enables interaction with humans but also exhibits superior performance in navigation and obstacle avoidance compared to state-of-the-art DRL algorithms. We believe this innovative paradigm opens up new avenues for effective and socially aware human-robot interactions in dynamic environments. Videos are available at https://hsacllm.github.io/.
I INTRODUCTION
Robot navigation is a crucial research domain that enables a robot to move within an environment while avoiding collisions and reaching specified destinations [1]. It has found applications in many fields, such as autonomous vehicles [2], drones [3], warehouse robots [4], etc. However, traditional robot navigation systems often prioritize efficiency and obstacle avoidance, but lack a detailed understanding of social cues and human intentions, which are critical for harmonious coexistence in shared spaces.
With the remarkable advancements in the realm of robotics in recent years, there is a growing emphasis on endowing robots with the capability to navigate and interact within intricate real-world environments. Socially-aware navigation [5] has emerged as a pivotal research area, aiming to empower robots with seamless navigational skills while respecting social norms and engaging in natural human interactions. Beyond conventional navigational abilities, robots are expected to comprehend human social norms, predict human behavior, and prevent actions that might lead to discomfort or misunderstandings. For instance, robots with socially-aware navigation capabilities can be deployed for navigation within international airports. These robots harness sophisticated language comprehension abilities and a deep understanding of social norms, enabling them to engage with passengers from diverse cultural backgrounds in natural language, while observing appropriate interpersonal distances.
Figure 1: Strategies used to avoid incoming collision with pedestrians. Left: Detour for a longer route to prevent collision [6]. Middle: Beeping to alert pedestrians and create space for the robot [7]. Right: Interactive voice interaction, enabling natural language conversations between robots and pedestrians (our approach).
Earlier studies typically aimed at modeling or predicting human motion, allowing robots to devise safe paths. Works such as the Social Force Model (SFM) [8], Social LSTM [9], and Social GAN [10] focused on predicting human motion trends, enabling robots to map out safer trajectories. Additionally, some research recognized human gestures or gaze directions to foresee pedestrian motion and adapt navigational actions accordingly [11]. However, these methods still primarily concentrated on how robots could evade pedestrians, often leading to longer detour distances or even collisions due to inaccurate pedestrian motion predictions. To facilitate meaningful interactions between robots and nearby individuals, recent endeavors have utilized “Beeping”, “Voice noticing”, and “Nudge” [7, 12, 13] as alert mechanisms for pedestrians. However, these methods, offering only simplistic, one-way alerts from robots, not only fall short of enabling a truly comfortable and interactive experience but also risk ambiguity and misinterpretation, particularly in noisy or crowded settings where the alert’s intent or origin may be unclear. Furthermore, repetitive or loud alerts can become intrusive or annoying, especially in environments where quiet or decorum is valued, thus adversely affecting user experience and the acceptance of robot assistance.
In this paper, we propose the Hybrid Soft Actor-Critic with Large Language Model (HSAC-LLM), a pioneering model that integrates deep reinforcement learning (DRL) with large language models (LLM) for socially-aware robot navigation, achieving natural language conversations with humans along the path. Fig.1 illustrates our method of robot-human interaction through natural language conversations, comparing it with two other existing types of interaction approaches during navigation: non-interaction strategies (Detour[6]) and unidirectional voice interaction (Beep [7]). It is evident that our method enables robots to understand human intentions more intelligently and clearly, especially in complex environments. Specifically, we first process the raw environmental data through PreNet networks to capture the current state, among which the voice state input is derived from conversations that the robot initiates, either actively or passively, with pedestrians upon detecting a potential collision. These states are subsequently input into our HSAC-LLM model, which simultaneously models the distributions of the robot’s continuous actions (linear and angular speeds) and discrete actions (interaction codes). Moreover, the robot communicates its final actions to pedestrians by generating corresponding voice notifications using the LLM model. Experiments conducted in 2D simulations, the Gazebo environment, and real-world settings demonstrate the superior performance of the proposed HSAC-LLM model over contemporary state-of-the-art DRL algorithms. This innovative paradigm in robot navigation paves the way for efficient and socially-aware human-robot interactions in dynamic settings. It not only enhances the safety and efficiency of robot navigation but also fosters harmonious coexistence between humans and robots in intricate social scenarios.
II RELATED WORK
II-A Robot Navigation
Safely reaching a destination while avoiding collisions with obstacles has always been a fundamental challenge in the field of mobile robotics. Traditional approaches have involved the integration of Simultaneous Localization and Mapping (SLAM) [14]. These approaches [15, 16, 17] typically require the creation of highly precise environment maps, followed by the use of path-planning algorithms to achieve accurate navigation. However, they often face limitations when applied in dynamic environments that demand significant computational resources for real-time path replanning [18]. Recently, by combining the robust representational capabilities of deep neural networks with reinforcement learning, DRL approaches [7, 19, 20, 21] have been employed in robot navigation, yielding encouraging results in dynamic environments with raw sensor inputs.
Figure 2: Illustration of a typical environment and the architecture of the proposed HSAC-LLM model.
II-B Human-Robot Interaction in Robot Navigation
Interacting with humans is unavoidable for evolving robots in roles like food delivery [22], tour guiding [23], and healthcare services [24]. Ensuring comfort, naturalness, and sociability is crucial in human-aware navigation for wheeled robots [25]. Early research usually aimed to model or predict human movements, enabling safer path planning. Pioneering works such as the Social Force Model (SFM)[8], Social LSTM[9], Social GAN [10], and MSA3C [26] focused on predicting human trends for safer paths. Subsequent research employed computer vision to recognize gestures like hand signals or gaze direction [11], but often prioritized navigation over interaction, resulting in longer detours or immobilization when avoidance from humans was needed.
Recent algorithms enable proactive robot-human interactions during the navigation process. Voice noticing like “Beeping” has been used as a popular interaction method to notice surrounding pedestrians and clear the path for the robot. “Beeping” yields a higher success rate and lower collision rate with pedestrians in the indoor navigation environment [7], and also shows it superior in decreasing collision rate in the elevator entering task [12]. Moreover, gestures like “nudging” have also been applied in the robot navigation task and show an enhanced success rate in the navigation task [13]. However, these approaches rely on the robot’s initiative, and assertive actions like beeping or nudging may not consistently offer a comfortable and natural interaction experience.
II-C Discrete-Continuous Hybrid Action Space
Enabling robots to control continuous velocity while providing discrete interaction options for engaging with humans presents the challenge of managing a Discrete-Continuous hybrid action space. Although DRL algorithms with hybrid action space have been validated on other robotic tasks like robotic arm-related tasks [27], this hybrid action space poses difficulties in adapting to the majority of DRL algorithms that have proven effective in robot navigation tasks. Previous approaches have tackled this issue through various methods, such as categorizing the continuous velocity action space, which discretizes the entire action space [7], or employing Reinforcement Learning to select interaction options [28]. However, the latter approach often requires extensive human-labeled data for training and may introduce subjective labeling issues. Hybrid action spaces are also common in video games, including adventure and racing games. In these games, players smoothly control characters with continuous movements while occasionally making discrete decisions like jumping or braking. DRL approaches like Multi-Pass deep Q-Networks [29] and Hybrid Soft Actor-Critic [30] have been developed and validated in various hybrid action spaces Gym games, laying the foundation for our approach to human-robot interaction tasks in robot navigation scenarios.
III Methods
III-A Overview
We introduce the Hybrid Soft Actor-Critic with Large Language Model (HSAC-LLM), a novel paradigm that combines deep reinforcement learning (DRL) with Large Language Models (LLMs) to facilitate bidirectional natural language interactions between robots and pedestrians during navigation. This approach ultimately enables intelligent obstacle avoidance while providing a comfortable and interactive experience. Fig. 2 shows a typical environment and the architecture of the proposed HSAC-LLM model. The radar scans and pedestrian information are first input to an Img_PreNet network, obtaining the state $\bm{S}_{i}$. Concurrently, the robot’s position information is captured as state $\bm{S}_{p}$. When potential collision risks are detected, the robot either initiates or responds to vocal interactions with humans. These interactions are processed through a Lang_PreNet network to generate the conditional state $\bm{S}_{c}$. Combining these states, the robot derives its subsequent actions and velocities by feeding the aggregated states into the proposed HSAC-LLM. Crucially, this model leverages both state and condition in its decision-making, represented as $\pi(\bm{a}|[\bm{S}_{i},\bm{S}_{p}],\bm{S}_{c})$. Once an appropriate action is determined, the robot leverages Lang_GenNet to provide vocal feedback to pedestrians, facilitating real-time communication and enhancing safety during navigation in dynamic environments.
III-B PreNet
Fig. 2 illustrates our input data sources: a 180-degree 2D Laser scan grid map representing the robot’s spatial configuration, a three-channel pedestrian grid map indicating the nearby pedestrian’s location and speed, a 9-dimensional relative position vector for target position and orientation over three time steps, and voice messages from pedestrians. Rather than directly integrating these data into HSAC-LLM, we employ three preprocessing networks to convert image and voice data from the environment into state vectors.
Img_PreNet
Following [6], we initiate feature extraction using three convolutional layers and three max pooling layers on laser scan data and pedestrian information grid map, producing a 512-dimensional vector $\bm{S}_{i}$ through a fully connected layer.
Lang_PreNet
When the robot perceives a potential collision with a human, it engages in a vocal conversation, either initiating or responding. To better understand these voice information, we fed these messages into a Lang_PreNet network to produce state vector $\bm{S}_{c}$ for the robot policy condition.
State_PreNet
The image info vector $\bm{S}_{i}$, the relative position info vector $\bm{S}_{p}$ and condition code vector $\bm{S}_{c}$ are concatenated and passed into the State_PreNet which contains one fully connected layer with 512 units to form a 512-dimensional vector $S$, which will serve as the environment state for the HSAC-LLM model.
III-C HSAC-LLM
In this section, we introduce HSAC-LLM, a model that enables bidirectional communication between robots and pedestrians to coordinate trajectories and prevent collisions. It comprises two main components: the HSAC (Hybrid Soft Actor Critic) module and the Lang_GenNet module. HSAC models both continuous and discrete robot actions, while Lang_GenNet serves a complementary purpose, which interprets and conveys the discrete outputs as human-friendly information through vocal communication, ensuring a seamless and safe co-navigation experience.
HSAC Module
During robot navigation, behavior involves both continuous and discrete elements. We categorize the robot’s motion controls, angular and linear velocities, as continuous actions, while post-interaction decisions are discrete actions. The standard SAC (Soft Actor-Critic) approach was originally introduced to model continuous actions for robotic tasks. Inspired by  [30], we propose a Hybrid SAC model tailored specifically for our current navigation task, emphasizing human-robot interactions. Fig. 2 shows the architecture of the proposed HSAC-LLM model. Specifically, we first utilize the actor to generate a shared hidden state representation $\bm{h}$ that is used to produce both a discrete distribution $\bm{\pi}$ as well as the mean $\bm{\mu}$ and standard deviation $\bm{\sigma}$ of the continuous component. The discrete action $\bm{a}_{c}$ is sampled from $\bm{\pi}$ and while the continuous actions, including linear velocity $\bm{v}$ and angular velocity $\bm{\omega}$, are computed as according the following formulation:
$[\bm{v},\bm{\omega}]=tanh(\bm{\mu}+\epsilon\bm{\sigma})$ (1)
After obtaining the continuous and discrete actions, the critic network takes both the state $\bm{s}$ and the continuous action $\bm{\mu}$ and $\bm{\omega}$ as input and predicts the Q-values of all discrete actions in its output layer. And the HSAC’s policy update formula will be:
$\pi^{*}=\arg\min_{\pi}\mathbb{E}_{\bm{a}\sim\pi}[\alpha\log\pi(\bm{a}|[\bm{S}_{i},\bm{S}_{p}],\bm{S}_{c}))-Q(\bm{s},\bm{\mu},\bm{\omega},\bm{a}_{c})]$ (2)
Lang_GenNet Module
The Lang_GenNet module, predicated on the architecture of Large Language Models (LLMs), is meticulously crafted to transform the action vector $\bm{a}_{c}$, derived from the HSAC model’s output, into coherent natural language interactions with pedestrians. In this study, the module is instantiated leveraging the capabilities of the GPT-4 model, showcasing a sophisticated integration of state-of-the-art language understanding and generation techniques.
III-D Reward Function
We define the reward for the robot at timestep $t$ as $r^{t}$, which is composed of four components: $r^{t}_{goal}$, $r^{t}_{shaping}$, $r^{t}_{step}$, and $r^{t}_{interact}$. For $r^{t}_{goal}$, the robot is awarded a positive reward $r_{\text{arrive}}$, if it successfully reaches its destination without its minimum distance to pedestrian $d^{t}_{min}$ falling below the safety threshold $d_{safe}$ at any point during its journey. However, if the robot reaches its goal but at any moment the minimum distance to a pedestrian $d^{t}_{min}$ drops below the safety threshold $d_{safe}$, a negative penalty proportional to the infraction is applied to $r_{\text{arrive}}$. In the event of a collision, $r^{t}_{goal}$ incurs a negative penalty, denoted as $r_{\text{collide}}$. In all other cases, $r^{t}_{goal}$ is set to 0. We formalize this as follows:
$r^{t}_{goal}={}\begin{cases}r_{arrive}&\quad\text{if }arrive,\\ r_{arrive}-\epsilon_{1}(d_{safe}-d^{t}_{min})&\quad\text{if }d^{t}_{min}<d_{safe},\\ r_{collide}&\quad\text{if }collision,\\ 0&\quad otherwise,\end{cases}$ (3)
where $\epsilon_{1}$ is a hyper-parameter. And $r^{t}_{\text{shaping}}$, motivates the robot to advance towards its target position by leveraging its current position $p^{t}$ and the target position $p_{g}$. This is defined as follows:
$r^{t}_{shaping}={}\epsilon_{2}(||p^{t-1}-p_{g}||-||p^{t}-p_{g}||)$ (4)
where $\epsilon_{2}$ is a hyper-parameter. Moreover, we employ the number of steps taken as a negative penalty, $r^{t}_{step}$, which encourages the robot to select shorter paths, thereby increasing its movement efficiency.
In addition, we introduced an interaction reward, $r^{t}_{interact}$, designed to encourage proper interaction between the robot and humans. When the robot initiates a dialogue request, $R_{r}$, and this request is in alignment with the current pedestrian trajectory, a reward of $r^{t}_{req}$ is granted. Conversely, if it is the pedestrian who initiates the dialogue request, $R_{p}$, and the robot’s path planning generated by the model aligns with the pedestrian’s request, a reward of $r^{t}_{res}$ is bestowed. In all other instances, $r^{t}_{interact}$ is set to 0. The formulation is as follows:
$r^{t}_{interact}={}\begin{cases}r^{t}_{req}&\quad\text{if }R_{r}\text{ is not none}\\ r^{t}_{res}&\quad\text{if }R_{p}\text{ is not none}\\ 0&\quad otherwise.\\ \end{cases}$ (5)
Hence, the final reward $r^{t}$ can be calculated as:
$r^{t}=r^{t}_{goal}+r^{t}_{shaping}+r^{t}_{step}+r^{t}_{interact}$ (6)
IV Experiments
Our algorithm is trained within a 2D simulator, utilizing the OpenCV framework as detailed in [31] and [6]. Subsequently, we evaluate its performance across three distinct environments: the 2D simulation environment, the Gazebo simulation environment, and the real-world environment.
Figure 3: Visualization of navigation trajectories in the Square, Hallway, and Crosswalk scenarios for baseline models (top) and our HSAC-LLM model (bottom).
IV-A Experiments in 2D simulator
IV-A1 Experimental Setting
In the 2D simulator, all pedestrians were controlled by a customized version of the Reciprocal Velocity Obstacles (RVO) algorithm [32], which enabled them to respond to the robot’s requests to dodge by performing stop and margin actions. We refer to this augmented model as Interactive Reciprocal Velocity Obstacles (IRVO). Our experimental focus was on scenarios involving a single pedestrian, closely mirroring real-world conditions for voice-based human-robot interactions. For all algorithms, the action spaces for linear and angular velocities were set to $v\in[0,0.6]$ m/s and $\omega\in[-0.9,0.9]$ rad/s, respectively; the velocity of the pedestrian was fixed at 0.5 m/s. Equation 6 served as the reward function for the baseline methods. The GPT-4 model from OpenAI [33] was selected as the LLM for the HSAC-LLM model. We employed a four-dimensional discrete action space for action codes, denoted as $\bm{a}_{c}\in$ none, stop, margin-right, margin-left, where “none” indicates no interaction, “stop” signals the pedestrian to halt, and “margin” commands the pedestrian to move to the left or right for creating space.
Considering the complexity of real-world navigation environments, we designed three distinct scenes within the simulator: Square, Hallway, and Crosswalk, as illustrated in Fig. LABEL:fig:results. To simulate varied navigation challenges, the starting and ending positions of the robot and pedestrian within each scene adhere to specific patterns:
  • Square: The robot and pedestrian are placed at random locations within the map.
  • Hallway: Both the robot and the pedestrian commence from random starting points at opposite ends of the Hallway, to traverse to the other side.
  • Crosswalk: The robot and pedestrian initiate their paths from random positions at the ends of differing corridors, aiming to reach the opposite side.
The primary goal of the robot is to navigate to its destination, prioritizing pedestrian safety and collision avoidance. The parameters of the reward function were set as follows: $r_{arrive}=500$, $r_{collide}=-500$, $r_{step}=-5$, $d_{safe}=1.0$, $\epsilon_{1}=-50$ and $\epsilon_{2}=200$.
IV-A2 Metrics
Following the setting of [34], we have selected the following metrics for evaluation of the navigation performance of our methods:
  • Success: The ratio of the number of times the robot safely reaches the target to the total number of attempts.
  • Collision: The ratio of the number of collisions with walls or pedestrians to the total number of attempts.
  • Step: The average number of steps taken by the robot to successfully reach the target.
  • Training Time: The time taken to train the model.
  • Human Collisions: The ratio of the number of collisions with humans to the total number of attempts.
  • Minimum Distance to Human: The shortest distance between the robot and human during successful navigation to the target.
  • Comfort Space Compliance: The ratio of instances in which the minimum distance to a human exceeds the comfort distance threshold to the total number of attempts.
IV-A3 Baselines
We incorporated state-of-the-art DRL algorithms tailored for continuous action spaces as our baseline methods, including Advantage Actor Critic (A2C) [35], Proximal Policy Optimization(PPO)[36], and Twin Delayed Deep Deterministic Policy Gradients(TD3)[37].
IV-A4 Results
All models were trained on an i7-10750H CPU and an NVIDIA RTX 2070 GPU for 128,000 steps. The results of 500 tests in each scene are presented in Table I, and the trajectories of various agents are depicted in Fig. 3. Throughout testing, a reverse-request probability of $p_{r}=0.5$ was employed, ensuring an equal chance for both the robot and the pedestrian to initiate an avoidance request.
TABLE I: Performance of our HSAC-LLM model compared with different baseline models in three scenarios. MDH stands for Minimum Distance to Human, and CSC denotes Comfort Space Compliance
[…]
Square A2C 0.69 ± 0.045 0.28 ± 0.030 27.38 0.9 0.28 0.25 0.81
TD3 0.85 ± 0.009 0.14 ± 0.008 21.21 2.6 0.14 0.28 0.80
PPO 0.58 ± 0.028 0.42 ± 0.023 37.81 8.0 0.39 0.23 0.82
HSAC-LLM 0.91 ± 0.022 0.08 ± 0.022 19.96 1.6 0.07 0.31 0.85
Hallway A2C 0.28 ± 0.038 0.59 ± 0.096 58.85 0.9 0.32 0.42 0.91
TD3 0.77 ± 0.044 0.20 ± 0.035 41.12 5.0 0.21 0.30 0.89
PPO 0.63 ± 0.037 0.37 ± 0.037 30.30 6.6 0.08 0.50 0.91
HSAC-LLM 0.93 ± 0.013 0.07 ± 0.013 27.10 1.9 0.06 0.32 0.93
Crosswalk A2C 0.14 ± 0.016 0.77 ± 0.029 43.83 1.1 0.58 1.28 0.99
TD3 0.85 ± 0.053 0.14 ± 0.044 41.67 3.0 0.14 0.33 0.90
PPO 0.46 ± 0.065 0.53 ± 0.065 33.33 10.8 0.53 0.54 0.91
HSAC-LLM 0.91 ± 0.022 0.09 ± 0.022 32.80 2.0 0.08 0.40 0.92
From Table I, we observe that our HSAC-LLM algorithm outperforms the state-of-the-art DRL algorithms in all three scenarios, achieving the highest success rates, the lowest collision rates, and the shortest step counts. Specifically, in the Square scenario, we achieved a success rate of 0.91, a collision rate of 0.08, a minimum distance to humans of 0.31, and a Comfort Space Compliance of 0.85. The significantly lower collision rate compared to other methods highlights our model’s ability to greatly reduce collisions through natural language-based interactions. In the more complex Hallway and Crosswalk scenarios, where collisions between humans and robots are more likely, our model still maintains collision rates of only 0.07 and 0.09, respectively. This is in contrast to other methods, which rely solely on unidirectional detouring to avoid collisions, a tactic that often results in collisions in slightly more complex environments due to insufficient evasion time. However, our HSAC-LLM model flexibly engages in natural language-based communication with pedestrians before potential collisions occur, fully understanding their intentions and eliminating ambiguities in communication, thus effectively preventing any impending collision.
To more intuitively showcase navigational performance, we randomly selected examples from three scenarios and illustrated the navigation trajectories of both baseline models and our model in Fig. 3. We observed that the navigation paths of these baseline models are considerably circuitous, with the PPO model in particular experiencing collisions with pedestrians in both the Square and Crosswalk scenarios. In contrast, our model facilitates bidirectional natural language dialogue between the robot and pedestrians, enabling the dynamic adjustment of passage priorities based on the specific context. Consequently, it efficiently reaches the destination via shorter paths while avoiding collisions.
IV-B Experiments in Gazebo and Real-World
Figure 4: Interactions between robot and pedestrian in the “Crosswalk” scene with the Gazebo environment. Figures (a) and (b) respectively illustrate explicit and implicit requests from pedestrians, demonstrating our model’s capability to comprehend natural language and respond accordingly.
To enhance the practicality of our algorithm, we crafted realistic 3D scenes within the Gazebo simulation environment. In these scenarios, pedestrian speeds are no longer fixed at 0.5 m/s but are estimated through a Leg Tracking framework [38], enabling the robot to dynamically decide the appropriate distance for initiating or responding to pedestrian dialogues. Pedestrian voice inputs and robot responses are simulated using the computer’s microphone and speakers. Extensive evaluations across three scenes revealed that the robot consistently demonstrates high responsiveness, navigating according to pedestrian preferences, either proactively or reactively. An example shown in Fig. 4 features a robot in the crosswalk scene obeying a pedestrian’s request to halt, allowing the pedestrian to pass safely. Notably, our model is adept at comprehending both explicit requests, like “Could you let me pass first?” and implicit ones, such as “I’m in a hurry,” thereby responding appropriately. This illustrates our model’s capability to understand the deeper nuances of human natural language, beyond mere literal interpretation. By deeply understanding human needs, our model can plan reasonable navigational routes to fulfill those needs. More results from the Gazebo simulation environment are available through videos on our website [^1] https://hsacllm.github.io/.
Furthermore, we conducted real-world environment tests to evaluate the navigational capabilities of our model using the TurtleBot3 robot. We opted for the Burger model, as depicted in Fig. 5, equipped with an RPLIDAR A1 2D lidar, an Intel D435i depth camera, and an NUC11PHKi7C computer. These tests revealed that, even in practical settings, our model continued to demonstrate commendable navigational performance. For example, when pedestrians explicitly expressed their desire to pass on the left, the robot strategically planned to move to the right and successfully executed the necessary maneuvers. This further highlights the effectiveness of our model in interacting with humans and implementing sensible path planning and actions in real-world scenarios. The video for the real-world environment can also be found on our website.
Figure 5: The TurtleBot3 Burger robot and environment scenarios for real-world experiments.
IV-C Ablation Study
IV-C1 Discrete Action Space Dimension
To evaluate the impact of the dimension of the discrete action space, we compared the navigational performance of the HSAC-LLM model employing both 2-D and 4-D discrete action spaces. Specifically, we utilized a common 2-D discrete action space, defined as $\bm{a}_{c}\in{\text{none},\text{stop}}$. Table II presents the navigational outcomes using the HSAC-LLM model with these two configurations. As indicated in the table, despite requiring a slightly longer training duration, the HSAC-LLM model utilizing a 4-D discrete action space surpasses the performance of its 2-D counterpart across all metrics, except for training time. This enhancement is particularly pronounced in scenarios where the robot and pedestrian are in close proximity and need to adjust their positions to avoid collisions.
TABLE II: Performance of HSAC-LLM trained in 2-D and 4-D discrete action space, best values bolded.
[…]
Square HSAC-LLM(2-D) HSAC-LLM(4-D) 0.90 $\pm$ 0.029 0.91 $\pm$ 0.022 0.10 $\pm$ 0.029 0.08 $\pm$ 0.022 20.84 19.96 1.45 1.60
Hallway HSAC-LLM(2-D) HSAC-LLM(4-D) 0.92 $\pm$ 0.036 0.93 $\pm$ 0.013 0.08 $\pm$ 0.036 0.07 $\pm$ 0.013 29.56 27.10 1.67 1.89
Crosswalk HSAC-LLM(2-D) HSAC-LLM(4-D) 0.87 $\pm$ 0.026 0.91 $\pm$ 0.022 0.13 $\pm$ 0.026 0.09 $\pm$ 0.022 35.22 32.80 1.78 2.04
IV-C2 Effectiveness of LLM module
In this section, we validate the effectiveness of the LLM component within the HSAC-LLM model. To this end, we conducted a comparative experiment by removing the LLM module responsible for natural language bidirectional interaction with pedestrians from the HSAC-LLM model. Specifically, when the robot detected a potential collision risk, it merely issued an auditory warning to pedestrians without the capability to receive vocal feedback from them. Table III presents a comparison between the HSAC model, with the LLM module removed, and the original HSAC-LLM model across three scenarios. The results from the table show that with the incorporation of the LLM module, there was an increase in success rates and a decrease in both collision rates and the number of steps to reach the target across all three scenarios. Notably, the improvement in collision rates was more significant, especially in the Square and Crosswalk scenarios, with reductions of 42.86% and 30.77%, respectively. This demonstrates that the integration of LLM enables the robot to better understand natural language and communicate with pedestrians, thereby eliminating collisions that could arise from inadequate communication. The inclusion of the LLM module not only markedly improves navigational safety by minimizing potential collisions but also significantly enhances the robot’s capability in social navigation. By enabling more nuanced natural language communication with pedestrians, this development highlights the essential role of advanced language processing in enriching human-robot interactions within dynamic environments. This progress in social navigation illustrates the substantial potential of integrating sophisticated language models into robotic systems, paving the way for more intuitive and effective human-robot collaborations, especially in complex social settings.
TABLE III: Performance of HSAC-LLM trained in the 1-way and 2-way interaction scenarios, best values bolded.
[…]
Square HSAC HSAC-LLM 0.86 $\pm$ 0.023 0.91 $\pm$ 0.022 0.14 $\pm$ 0.023 0.08 $\pm$ 0.022 20.02 19.96 0.14 0.07
Hallway HSAC HSAC-LLM 0.91 $\pm$ 0.018 0.93 $\pm$ 0.013 0.09 $\pm$ 0.018 0.07 $\pm$ 0.013 29.00 27.10 0.09 0.06
Crosswalk HSAC HSAC-LLM 0.87 $\pm$ 0.026 0.91 $\pm$ 0.022 0.13 $\pm$ 0.026 0.09 $\pm$ 0.022 34.78 32.80 0.13 0.08
V CONCLUSIONS
This paper introduces HSAC-LLM, a novel approach to socially-aware navigation that combines deep reinforcement learning with large language models. This approach enables bidirectional interactions based on natural language between robots and pedestrians. Experiments in 2D simulations, the Gazebo environment, and real-world settings further demonstrate the superior performance of our model in navigation and obstacle avoidance. We believe this innovative paradigm not only enhances the safety and efficiency of robot navigation but also facilitates harmonious coexistence between humans and robots in complex social environments.

### v2

@@
-Enhancing Socially-Aware Robot Navigation through Bidirectional Natural Language Conversation
+Socially-Aware Robot Navigation Enhanced by Bidirectional Natural Language Conversations Using Large Language Models
 Abstract
-Robot navigation is an important research field with applications in various domains.
-However, traditional approaches often prioritize efficiency and obstacle avoidance, neglecting a nuanced understanding of human behavior or intent in shared spaces.
-With the rise of service robots, there’s an increasing emphasis on endowing robots with the capability to navigate and interact in complex real-world environments.
-Socially aware navigation has recently become a key research area.
-However, existing work either predicts pedestrian movements or simply emits alert signals to pedestrians, falling short of facilitating genuine interactions between humans and robots.
-In this paper, we introduce the Hybrid Soft Actor-Critic with Large Language Model (HSAC-LLM), an innovative model designed for socially-aware navigation in robots.
-This model seamlessly integrates deep reinforcement learning with large language models, enabling it to predict both continuous and discrete actions for navigation.
-Notably, HSAC-LLM facilitates bidirectional interaction based on natural language with pedestrian models.
-When a potential collision with pedestrians is detected, the robot can initiate or respond to communications with pedestrians, obtaining and executing subsequent avoidance strategies.
-Experimental results in 2D simulation, the Gazebo environment, and the real-world environment demonstrate that HSAC-LLM not only efficiently enables interaction with humans but also exhibits superior performance in navigation and obstacle avoidance compared to state-of-the-art DRL algorithms.
-We believe this innovative paradigm opens up new avenues for effective and socially aware human-robot interactions in dynamic environments.
+Robot navigation is crucial across various domains, yet traditional methods focus on efficiency and obstacle avoidance, often overlooking human behavior in shared spaces.
+With the rise of service robots, socially aware navigation has gained prominence.
+However, existing approaches primarily predict pedestrian movements or issue alerts, lacking true human-robot interaction.
+We introduce Hybrid Soft Actor-Critic with Large Language Model (HSAC-LLM), a novel framework for socially aware navigation.
+By integrating deep reinforcement learning with large language models, HSAC-LLM enables bidirectional natural language interactions, predicting both continuous and discrete navigation actions.
+When potential collisions arise, the robot proactively communicates with pedestrians to determine avoidance strategies.
+Experiments in 2D simulation, Gazebo, and real-world environments demonstrate that HSAC-LLM outperforms state-of-the-art DRL methods in interaction, navigation, and obstacle avoidance.
+This paradigm advances effective human-robot interactions in dynamic settings.
 Videos are available at https://hsacllm.github.io/.
 I INTRODUCTION
-Robot navigation is a crucial research domain that enables a robot to move within an environment while avoiding collisions and reaching specified destinations [1].
-It has found applications in many fields, such as autonomous vehicles [2], drones [3], warehouse robots [4], etc.
-However, traditional robot navigation systems often prioritize efficiency and obstacle avoidance, but lack a detailed understanding of social cues and human intentions, which are critical for harmonious coexistence in shared spaces.
-With the remarkable advancements in the realm of robotics in recent years, there is a growing emphasis on endowing robots with the capability to navigate and interact within intricate real-world environments.
-Socially-aware navigation [5] has emerged as a pivotal research area, aiming to empower robots with seamless navigational skills while respecting social norms and engaging in natural human interactions.
-Beyond conventional navigational abilities, robots are expected to comprehend human social norms, predict human behavior, and prevent actions that might lead to discomfort or misunderstandings.
-For instance, robots with socially-aware navigation capabilities can be deployed for navigation within international airports.
-These robots harness sophisticated language comprehension abilities and a deep understanding of social norms, enabling them to engage with passengers from diverse cultural backgrounds in natural language, while observing appropriate interpersonal distances.
+Robot navigation is essential for enabling movement while avoiding collisions across applications such as autonomous vehicles [1], drones [2], and robots [3, 4].
+However, traditional approaches prioritize efficiency and obstacle avoidance while neglecting social cues and human intentions, crucial for coexistence in shared spaces.
+With advances in robotics, socially-aware navigation [5] has emerged as a key area, equipping robots with the ability to navigate while respecting social norms and engaging in natural interactions.
+Beyond safe movement, robots must predict human behavior and adhere to social etiquette.
+For instance, airport navigation robots leverage language comprehension and social awareness to interact naturally with passengers from diverse backgrounds.
+Earlier methods focused on human motion prediction for safer path planning, using models like Social Force Model (SFM)[6], Social LSTM[7], and Social GAN [8].
+Gesture and gaze recognition techniques [9] further improved predictions but often led to unnecessary detours or collisions.
+Recent approaches introduce pedestrian alerts through ”Beeping,” ”Voice Noticing,” and ”Nudging” [10, 11, 12].
+However, these one-way signals can be ambiguous, ineffective in noisy environments, and intrusive, diminishing user experience and robot acceptance.
+In this paper, we propose the Hybrid Soft Actor-Critic with Large Language Model (HSAC-LLM), integrating deep reinforcement learning with large language models for socially-aware robot navigation with natural language capabilities.
+Fig.1 compares our approach with existing methods: non-interaction strategies (Detour[13]) and unidirectional voice interaction (Beep [10]).
+Our method enables robots to better understand human intentions in complex environments.
+We process environmental data through PreNet networks to capture the current state, including voice input from conversations initiated when potential collisions are detected.
+Our HSAC-LLM model simultaneously predicts continuous actions (linear/angular speeds) and discrete actions (interaction codes), while generating appropriate voice notifications via the LLM.
+Experiments in 2D simulations, Gazebo environments, and real-world settings demonstrate HSAC-LLM’s superior performance over state-of-the-art DRL algorithms.
+This approach enhances navigation safety and efficiency while fostering harmonious human-robot coexistence in social scenarios.
+The main contributions are as follows:
 Figure 1: Strategies used to avoid incoming collision with pedestrians.
@@
 Right: Interactive voice interaction, enabling natural language conversations between robots and pedestrians (our approach).
-Earlier studies typically aimed at modeling or predicting human motion, allowing robots to devise safe paths.
-Works such as the Social Force Model (SFM) [8], Social LSTM [9], and Social GAN [10] focused on predicting human motion trends, enabling robots to map out safer trajectories.
-Additionally, some research recognized human gestures or gaze directions to foresee pedestrian motion and adapt navigational actions accordingly [11].
-However, these methods still primarily concentrated on how robots could evade pedestrians, often leading to longer detour distances or even collisions due to inaccurate pedestrian motion predictions.
-To facilitate meaningful interactions between robots and nearby individuals, recent endeavors have utilized “Beeping”, “Voice noticing”, and “Nudge” [7, 12, 13] as alert mechanisms for pedestrians.
-However, these methods, offering only simplistic, one-way alerts from robots, not only fall short of enabling a truly comfortable and interactive experience but also risk ambiguity and misinterpretation, particularly in noisy or crowded settings where the alert’s intent or origin may be unclear.
-Furthermore, repetitive or loud alerts can become intrusive or annoying, especially in environments where quiet or decorum is valued, thus adversely affecting user experience and the acceptance of robot assistance.
-In this paper, we propose the Hybrid Soft Actor-Critic with Large Language Model (HSAC-LLM), a pioneering model that integrates deep reinforcement learning (DRL) with large language models (LLM) for socially-aware robot navigation, achieving natural language conversations with humans along the path.
-Fig.1 illustrates our method of robot-human interaction through natural language conversations, comparing it with two other existing types of interaction approaches during navigation: non-interaction strategies (Detour[6]) and unidirectional voice interaction (Beep [7]).
-It is evident that our method enables robots to understand human intentions more intelligently and clearly, especially in complex environments.
-Specifically, we first process the raw environmental data through PreNet networks to capture the current state, among which the voice state input is derived from conversations that the robot initiates, either actively or passively, with pedestrians upon detecting a potential collision.
-These states are subsequently input into our HSAC-LLM model, which simultaneously models the distributions of the robot’s continuous actions (linear and angular speeds) and discrete actions (interaction codes).
-Moreover, the robot communicates its final actions to pedestrians by generating corresponding voice notifications using the LLM model.
-Experiments conducted in 2D simulations, the Gazebo environment, and real-world settings demonstrate the superior performance of the proposed HSAC-LLM model over contemporary state-of-the-art DRL algorithms.
-This innovative paradigm in robot navigation paves the way for efficient and socially-aware human-robot interactions in dynamic settings.
-It not only enhances the safety and efficiency of robot navigation but also fosters harmonious coexistence between humans and robots in intricate social scenarios.
+  • We explore the integration of socially-aware navigation and natural language interaction, addressing the challenge of enabling robots to navigate complex and dynamic environments while respecting social norms.
+  • We develop HSAC-LLM, a unified framework combining deep reinforcement learning and large language models that empower robots to understand and respond to human intentions through natural conversation.
+  • We introduce comprehensive evaluation metrics for socially-aware navigation and rigorously assess its performance from three aspects: safety, efficiency, and human comfort.
+  • We validate our approach through experiments in simulated and real-world environments, demonstrating superior performance over existing methods.
 II RELATED WORK
 II-A Robot Navigation
-Safely reaching a destination while avoiding collisions with obstacles has always been a fundamental challenge in the field of mobile robotics.
-Traditional approaches have involved the integration of Simultaneous Localization and Mapping (SLAM) [14].
-These approaches [15, 16, 17] typically require the creation of highly precise environment maps, followed by the use of path-planning algorithms to achieve accurate navigation.
-However, they often face limitations when applied in dynamic environments that demand significant computational resources for real-time path replanning [18].
-Recently, by combining the robust representational capabilities of deep neural networks with reinforcement learning, DRL approaches [7, 19, 20, 21] have been employed in robot navigation, yielding encouraging results in dynamic environments with raw sensor inputs.
+Safely reaching a destination while avoiding obstacles has long been a key challenge in mobile robotics.
+Early approaches relied on deterministic algorithms such as Simultaneous Localization and Mapping (SLAM) [14], which generated precise environment maps using sensor data like LiDAR and vision-based inputs [15].
+Path-planning methods, including A* and Dijkstra’s, were then applied to compute collision-free trajectories [16].
+However, these approaches struggled in dynamic environments due to high computational demands and limited adaptability [17].
+The emergence of Deep Reinforcement Learning (DRL) introduced a paradigm shift by leveraging raw sensor data to learn navigation policies without explicit mapping.
+DRL-enabled mapless navigation has demonstrated adaptability in dynamic environments with moving obstacles [18].
+Additionally, combining DRL with convolutional neural networks (CNNs) enhances feature extraction from sensor inputs, improving navigation in complex scenarios [19].
+Virtual-to-real transfer techniques further optimize DRL training, reducing real-world deployment costs [20].
 Figure 2: Illustration of a typical environment and the architecture of the proposed HSAC-LLM model.
 II-B Human-Robot Interaction in Robot Navigation
-Interacting with humans is unavoidable for evolving robots in roles like food delivery [22], tour guiding [23], and healthcare services [24].
-Ensuring comfort, naturalness, and sociability is crucial in human-aware navigation for wheeled robots [25].
-Early research usually aimed to model or predict human movements, enabling safer path planning.
-Pioneering works such as the Social Force Model (SFM)[8], Social LSTM[9], Social GAN [10], and MSA3C [26] focused on predicting human trends for safer paths.
-Subsequent research employed computer vision to recognize gestures like hand signals or gaze direction [11], but often prioritized navigation over interaction, resulting in longer detours or immobilization when avoidance from humans was needed.
-Recent algorithms enable proactive robot-human interactions during the navigation process.
-Voice noticing like “Beeping” has been used as a popular interaction method to notice surrounding pedestrians and clear the path for the robot.
-“Beeping” yields a higher success rate and lower collision rate with pedestrians in the indoor navigation environment [7], and also shows it superior in decreasing collision rate in the elevator entering task [12].
-Moreover, gestures like “nudging” have also been applied in the robot navigation task and show an enhanced success rate in the navigation task [13].
-However, these approaches rely on the robot’s initiative, and assertive actions like beeping or nudging may not consistently offer a comfortable and natural interaction experience.
+Robots increasingly interact with humans in roles like food delivery [21], tour guiding [22], and healthcare [23].
+Ensuring comfort, naturalness, and sociability is key for human-aware navigation [24].
+Early research focused on predicting human movements for safer path planning, employing models like the Social Force Model (SFM)[6], Social LSTM[7], Social GAN [8], and MSA3C [25].
+Computer vision techniques further enabled gesture recognition [9], but often prioritized navigation over interaction, leading to detours or immobilization.
+Recent methods integrate proactive robot-human interactions.
+Techniques like ”beeping” effectively alert pedestrians, improving success and reducing collisions in indoor navigation [10] and elevator entry tasks [11].
+Gestures like ”nudging” also enhance navigation performance [12].
+However, these methods rely on the robot’s initiative, and assertive actions may not always ensure a comfortable, natural interaction experience.
 II-C Discrete-Continuous Hybrid Action Space
-Enabling robots to control continuous velocity while providing discrete interaction options for engaging with humans presents the challenge of managing a Discrete-Continuous hybrid action space.
-Although DRL algorithms with hybrid action space have been validated on other robotic tasks like robotic arm-related tasks [27], this hybrid action space poses difficulties in adapting to the majority of DRL algorithms that have proven effective in robot navigation tasks.
-Previous approaches have tackled this issue through various methods, such as categorizing the continuous velocity action space, which discretizes the entire action space [7], or employing Reinforcement Learning to select interaction options [28].
-However, the latter approach often requires extensive human-labeled data for training and may introduce subjective labeling issues.
-Hybrid action spaces are also common in video games, including adventure and racing games.
-In these games, players smoothly control characters with continuous movements while occasionally making discrete decisions like jumping or braking.
-DRL approaches like Multi-Pass deep Q-Networks [29] and Hybrid Soft Actor-Critic [30] have been developed and validated in various hybrid action spaces Gym games, laying the foundation for our approach to human-robot interaction tasks in robot navigation scenarios.
+Managing a Discrete-Continuous hybrid action space is crucial for enabling robots to control continuous velocity while engaging in discrete human interactions.
+While DRL algorithms for hybrid action spaces have been validated in robotic arm tasks [26], adapting them to robot navigation remains challenging.
+Prior approaches have addressed this by discretizing velocity actions [10] or using Reinforcement Learning for interaction selection [27], but the latter often relies on extensive human-labeled data, introducing subjectivity.
+Hybrid action spaces are also prevalent in video games, where players control movements continuously while making discrete decisions like jumping or braking.
+DRL methods such as Multi-Pass deep Q-Networks [28] and Hybrid Soft Actor-Critic [29] have demonstrated success in Gym environments, providing a foundation for our approach to human-robot interaction in navigation tasks.
 III Methods
@@
 Crucially, this model leverages both state and condition in its decision-making, represented as $\pi(\bm{a}|[\bm{S}_{i},\bm{S}_{p}],\bm{S}_{c})$.
-Once an appropriate action is determined, the robot leverages Lang_GenNet to provide vocal feedback to pedestrians, facilitating real-time communication and enhancing safety during navigation in dynamic environments.
+Once an appropriate action is determined, the robot leverages Lang_GenNet to provide vocal feedback to pedestrians, facilitating real-time communication and enhancing safety during navigation.
 III-B PreNet
@@
 Equation 6 served as the reward function for the baseline methods.
-The GPT-4 model from OpenAI [33] was selected as the LLM for the HSAC-LLM model.
+The GPT-4 model from OpenAI was selected as the LLM for the HSAC-LLM model.
 We employed a four-dimensional discrete action space for action codes, denoted as $\bm{a}_{c}\in$ none, stop, margin-right, margin-left, where “none” indicates no interaction, “stop” signals the pedestrian to halt, and “margin” commands the pedestrian to move to the left or right for creating space.
 Considering the complexity of real-world navigation environments, we designed three distinct scenes within the simulator: Square, Hallway, and Crosswalk, as illustrated in Fig.
-LABEL:fig:results.
-To simulate varied navigation challenges, the starting and ending positions of the robot and pedestrian within each scene adhere to specific patterns:
-  • Square: The robot and pedestrian are placed at random locations within the map.
-  • Hallway: Both the robot and the pedestrian commence from random starting points at opposite ends of the Hallway, to traverse to the other side.
-  • Crosswalk: The robot and pedestrian initiate their paths from random positions at the ends of differing corridors, aiming to reach the opposite side.
+3.
+To simulate varied navigation challenges, the starting and ending positions of the robot and pedestrian within each scene adhere to specific patterns.
+In the Square scene, the robot and pedestrian are placed at random locations.
+For the Hallway scene, both agents start at opposite ends, moving toward each other.
+In the Crosswalk scenario, agents start at different corridor ends, crossing paths.
 The primary goal of the robot is to navigate to its destination, prioritizing pedestrian safety and collision avoidance.
@@
 IV-A2 Metrics
-Following the setting of [34], we have selected the following metrics for evaluation of the navigation performance of our methods:
-  • Success: The ratio of the number of times the robot safely reaches the target to the total number of attempts.
-  • Collision: The ratio of the number of collisions with walls or pedestrians to the total number of attempts.
-  • Step: The average number of steps taken by the robot to successfully reach the target.
-  • Training Time: The time taken to train the model.
-  • Human Collisions: The ratio of the number of collisions with humans to the total number of attempts.
-  • Minimum Distance to Human: The shortest distance between the robot and human during successful navigation to the target.
-  • Comfort Space Compliance: The ratio of instances in which the minimum distance to a human exceeds the comfort distance threshold to the total number of attempts.
+Following [32], we propose the following metrics to evaluate navigation performance:
+  • Success: Ratio of safe target arrivals to total attempts.
+  • Collision: Ratio of collisions to total attempts.
+  • Step: Average steps for successful navigation.
+  • Training Time: Model training duration.
+  • Human Collisions: Ratio of human collisions to total attempts.
+  • Minimum Distance to Human: Shortest robot-human distance during successful navigation.
+  • Comfort Space Compliance: Ratio of instances maintaining minimum comfort distance.
 IV-A3 Baselines
@@
 However, our HSAC-LLM model flexibly engages in natural language-based communication with pedestrians before potential collisions occur, fully understanding their intentions and eliminating ambiguities in communication, thus effectively preventing any impending collision.
+Figure 4: Interaction Scenarios in Hallway and Crosswalk Environments (Robot Proactive and Pedestrian Proactive cases).
 To more intuitively showcase navigational performance, we randomly selected examples from three scenarios and illustrated the navigation trajectories of both baseline models and our model in Fig.
@@
 Consequently, it efficiently reaches the destination via shorter paths while avoiding collisions.
-IV-B Experiments in Gazebo and Real-World
-Figure 4: Interactions between robot and pedestrian in the “Crosswalk” scene with the Gazebo environment.
-Figures (a) and (b) respectively illustrate explicit and implicit requests from pedestrians, demonstrating our model’s capability to comprehend natural language and respond accordingly.
-To enhance the practicality of our algorithm, we crafted realistic 3D scenes within the Gazebo simulation environment.
-In these scenarios, pedestrian speeds are no longer fixed at 0.5 m/s but are estimated through a Leg Tracking framework [38], enabling the robot to dynamically decide the appropriate distance for initiating or responding to pedestrian dialogues.
-Pedestrian voice inputs and robot responses are simulated using the computer’s microphone and speakers.
-Extensive evaluations across three scenes revealed that the robot consistently demonstrates high responsiveness, navigating according to pedestrian preferences, either proactively or reactively.
-An example shown in Fig.
-4 features a robot in the crosswalk scene obeying a pedestrian’s request to halt, allowing the pedestrian to pass safely.
-Notably, our model is adept at comprehending both explicit requests, like “Could you let me pass first?” and implicit ones, such as “I’m in a hurry,” thereby responding appropriately.
-This illustrates our model’s capability to understand the deeper nuances of human natural language, beyond mere literal interpretation.
-By deeply understanding human needs, our model can plan reasonable navigational routes to fulfill those needs.
-More results from the Gazebo simulation environment are available through videos on our website [^1] https://hsacllm.github.io/.
-Furthermore, we conducted real-world environment tests to evaluate the navigational capabilities of our model using the TurtleBot3 robot.
-We opted for the Burger model, as depicted in Fig.
-5, equipped with an RPLIDAR A1 2D lidar, an Intel D435i depth camera, and an NUC11PHKi7C computer.
-These tests revealed that, even in practical settings, our model continued to demonstrate commendable navigational performance.
-For example, when pedestrians explicitly expressed their desire to pass on the left, the robot strategically planned to move to the right and successfully executed the necessary maneuvers.
-This further highlights the effectiveness of our model in interacting with humans and implementing sensible path planning and actions in real-world scenarios.
-The video for the real-world environment can also be found on our website.
-Figure 5: The TurtleBot3 Burger robot and environment scenarios for real-world experiments.
-IV-C Ablation Study
-IV-C1 Discrete Action Space Dimension
-To evaluate the impact of the dimension of the discrete action space, we compared the navigational performance of the HSAC-LLM model employing both 2-D and 4-D discrete action spaces.
-Specifically, we utilized a common 2-D discrete action space, defined as $\bm{a}_{c}\in{\text{none},\text{stop}}$.
-Table II presents the navigational outcomes using the HSAC-LLM model with these two configurations.
-As indicated in the table, despite requiring a slightly longer training duration, the HSAC-LLM model utilizing a 4-D discrete action space surpasses the performance of its 2-D counterpart across all metrics, except for training time.
-This enhancement is particularly pronounced in scenarios where the robot and pedestrian are in close proximity and need to adjust their positions to avoid collisions.
+IV-B Experiments in Gazebo simulation
+To enhance the real-world applicability of our algorithm, we developed realistic 3D simulation scenarios using the Gazebo simulation environment.
+Unlike conventional setups where pedestrian speeds are fixed at 0.5 m/s, our approach incorporates a Leg Tracking framework [36] to dynamically estimate pedestrian speeds.
+This method enables the robot to adaptively determine the optimal distance for initiating or responding to pedestrian interactions, ensuring more flexible and natural movement planning.
+Importantly, the simulated human in the Gazebo environment is controlled by an actual person, who actively manages speed and direction to replicate realistic human behavior.
+In parallel, to simulate realistic and naturalistic human-robot interactions, we employed the computer’s microphone and speakers to facilitate pedestrian voice inputs and robotic responses.
+This setup allows the robot to process spoken commands or requests from pedestrians and generate appropriate verbal replies, creating a lifelike communication exchange.
+We conducted comprehensive evaluations across three distinct simulated environments to assess the robot’s performance under varying interaction scenarios.
+The results consistently demonstrate the robot’s high responsiveness and adaptability, effectively navigating according to pedestrian preferences.
+Specifically, the robot behaves proactively to avoid potential conflicts or reactively to accommodate pedestrian-initiated requests, depending on the situation.
+To intuitively illustrate the robot’s navigation performance, Fig.
+4 depicts two primary environments: Hallway and Crosswalk.
+Each environment contains two types of proactive behaviors: Hallway (Robot Proactive), Hallway (Pedestrian Proactive), Crosswalk (Robot Proactive), and Crosswalk (Pedestrian Proactive).
+For each scenario, we present five key stages of the interaction process: Before Encounter, Interaction Starts, Interaction Dialogue, Avoidance in Progress, and After Encounter.
+These results collectively demonstrate the versatility and robustness of our model in handling diverse interaction contexts, ensuring smooth and efficient navigation while adhering to pedestrian preferences.
+Notably, our model demonstrates the ability to successfully interpret both explicit commands and implicit cues.
+Explicit commands, such as “Could you let me pass first?”, are straightforward requests that the robot processes and responds to appropriately.
+However, the model also handles more subtle implicit cues, such as “I’m in a hurry,” with contextually appropriate behavior.
+For example, in the Crosswalk (Pedestrian Proactive) scenario, when the pedestrian states, “I’m in a hurry,” the robot responds with: “OK, I will stop to let you pass,” while simultaneously halting to enable the pedestrian to pass safely.
+This highlights the model’s ability to comprehend nuanced human language, extending beyond literal word interpretation to capture the underlying intent.
+By deeply understanding human preferences and needs, our model effectively plans and executes rational and context-aware navigational routes that align with those needs.
+This ability to process natural language, interpret human intent, and dynamically adjust navigation strategies underscores the robustness and practicality of our approach for real-world human-robot interactions.
+Figure 5: Real-world hallway experiments showing Robot Proactive (top) and Human Proactive (bottom) scenarios, demonstrates how natural-language communication enables effective collision avoidance between robot and pedestrian.
+IV-C Experiments in Real-World environment
+To evaluate our model’s practical efficacy, we implemented HSAC-LLM on a customized Hiwonder JetAuto Pro robot, as shown in Fig.
+5.
+This platform integrates an NVIDIA Jetson Nano (quad-core Cortex-A57 processor at 1.43 GHz with 128 Maxwell CUDA cores delivering 472 GFLOPS), an architecture well-suited for edge-based machine learning inference.
+The robot is equipped with an RPLIDAR A1 sensor, offering a 5–10 Hz scanning frequency, a 12-meter maximum range, and 1° angular resolution, enabling precise 2D environmental perception without additional visual sensors.
+The JetAuto Pro’s motorized chassis ensures reliable navigation, while the Jetson Nano’s energy efficiency supports real-time inference with minimal power consumption.
+Our real-world hallway experiments demonstrated two key interaction scenarios shown in Fig.
+5.
+Each scenario progresses through five phases: Before Encounter, Interaction Starts, Interaction Dialogue, Avoidance in Progress, and After Encounter.
+In the Robot Proactive scenario (top row), the robot requests ”Please move to your right!” and the human complies.
+In the Human Proactive scenario (bottom row), when warned ”Watch out on your right!”, the robot responds ”I’ll move to my left!” and adjusts accordingly.
+Results confirm that HSAC-LLM achieves robust real-time decision-making and collision avoidance using only a 2D LiDAR.
+This integration of natural language understanding with navigation validates our model’s feasibility for real-world applications requiring effective human-robot communication.
 TABLE II: Performance of HSAC-LLM trained in 2-D and 4-D discrete action space, best values bolded.
@@
 Crosswalk HSAC-LLM(2-D) HSAC-LLM(4-D) 0.87 $\pm$ 0.026 0.91 $\pm$ 0.022 0.13 $\pm$ 0.026 0.09 $\pm$ 0.022 35.22 32.80 1.78 2.04
-IV-C2 Effectiveness of LLM module
-In this section, we validate the effectiveness of the LLM component within the HSAC-LLM model.
-To this end, we conducted a comparative experiment by removing the LLM module responsible for natural language bidirectional interaction with pedestrians from the HSAC-LLM model.
-Specifically, when the robot detected a potential collision risk, it merely issued an auditory warning to pedestrians without the capability to receive vocal feedback from them.
-Table III presents a comparison between the HSAC model, with the LLM module removed, and the original HSAC-LLM model across three scenarios.
-The results from the table show that with the incorporation of the LLM module, there was an increase in success rates and a decrease in both collision rates and the number of steps to reach the target across all three scenarios.
-Notably, the improvement in collision rates was more significant, especially in the Square and Crosswalk scenarios, with reductions of 42.86% and 30.77%, respectively.
-This demonstrates that the integration of LLM enables the robot to better understand natural language and communicate with pedestrians, thereby eliminating collisions that could arise from inadequate communication.
-The inclusion of the LLM module not only markedly improves navigational safety by minimizing potential collisions but also significantly enhances the robot’s capability in social navigation.
-By enabling more nuanced natural language communication with pedestrians, this development highlights the essential role of advanced language processing in enriching human-robot interactions within dynamic environments.
-This progress in social navigation illustrates the substantial potential of integrating sophisticated language models into robotic systems, paving the way for more intuitive and effective human-robot collaborations, especially in complex social settings.
+IV-D Ablation Study
+IV-D1 Discrete Action Space Dimension
+To assess how discrete action space dimensionality affects navigation performance, we tested HSAC-LLM with both 2-D ($\bm{a}_{c}\in\{\text{none},\text{stop}\}$) and 4-D discrete action spaces across Square, Hallway, and Crosswalk environments.
+Table II presents the results based on Success Rate, Collision Rate, Steps Taken, and Training Time metrics.
+The 4-D discrete action space consistently outperformed the 2-D configuration across all environments.
+In the Square environment, success rate improved from 0.90 $\pm$ 0.029 to 0.91 $\pm$ 0.022, while collision rate decreased from 0.10 $\pm$ 0.029 to 0.08 $\pm$ 0.022.
+Similar improvements occurred in the Hallway scenario (success rate: 0.92 $\pm$ 0.036 to 0.93 $\pm$ 0.013; collision rate: 0.08 $\pm$ 0.036 to 0.07 $\pm$ 0.013).
+The Crosswalk environment showed the most significant improvement, with success rate increasing from 0.87 $\pm$ 0.026 to 0.91 $\pm$ 0.022 and collision rate decreasing from 0.13 $\pm$ 0.026 to 0.09 $\pm$ 0.022.
+While the 4-D configuration required slightly longer training times (e.g., increasing from 1.67 to 1.89 in Hallway and from 1.78 to 2.04 in Crosswalk), the performance gains justify this trade-off.
+These results demonstrate that expanding the discrete action space enhances HSAC-LLM’s ability to navigate complex environments safely and efficiently, particularly in scenarios requiring nuanced interactions with pedestrians.
 TABLE III: Performance of HSAC-LLM trained in the 1-way and 2-way interaction scenarios, best values bolded.
@@
 Crosswalk HSAC HSAC-LLM 0.87 $\pm$ 0.026 0.91 $\pm$ 0.022 0.13 $\pm$ 0.026 0.09 $\pm$ 0.022 34.78 32.80 0.13 0.08
+IV-D2 Effectiveness of LLM module
+To evaluate the LLM component’s effectiveness in HSAC-LLM, we conducted experiments comparing the full model against a version with the LLM module removed (HSAC).
+Without the LLM module, the robot could only issue warnings to pedestrians but couldn’t process their verbal feedback.
+Table III compares both models across three scenarios.
+Results show that incorporating the LLM module improved success rates and reduced collision rates and navigation steps in all scenarios.
+Collision rate improvements were particularly significant in Square and Crosswalk environments, with reductions of 42.86% and 30.77%, respectively.
+The LLM integration enables better natural language understanding and communication with pedestrians, preventing collisions caused by communication gaps.
+This enhancement not only improves navigational safety but also significantly advances the robot’s social navigation capabilities, demonstrating the value of sophisticated language processing in human-robot interactions within dynamic environments.
 V CONCLUSIONS
-This paper introduces HSAC-LLM, a novel approach to socially-aware navigation that combines deep reinforcement learning with large language models.
-This approach enables bidirectional interactions based on natural language between robots and pedestrians.
-Experiments in 2D simulations, the Gazebo environment, and real-world settings further demonstrate the superior performance of our model in navigation and obstacle avoidance.
-We believe this innovative paradigm not only enhances the safety and efficiency of robot navigation but also facilitates harmonious coexistence between humans and robots in complex social environments.
+This paper presented HSAC-LLM, a novel framework integrating deep reinforcement learning with large language models for socially-aware robot navigation.
+Our approach enables natural language interactions between robots and humans, enhancing the robot’s ability to understand human intentions in dynamic environments.
+HSAC-LLM models both continuous navigation actions and discrete conversation actions.
+Experiments in 2D simulations, Gazebo, and real-world settings demonstrate superior performance in safety, efficiency, and human comfort compared to state-of-the-art DRL methods.
+This work advances intelligent human-robot interactions in social environments.
+Future work will focus on improving language understanding, real-time adaptation to complex scenarios, and large-scale real-world deployment.